stormyfurr:

Simplify (−8.5)(−5)( −2) A. 850 B. 85 C. −85 D. −850

stormyfurr:

I need help

Imagine:

-5x-2=10 -8.5x10=?

stormyfurr:

-85

Imagine:

Yup! Nice.

stormyfurr:

Thanks

Imagine:

Your Welcome, Anytime! (: